Amir is 5 years younger than Alison, together their age totals to 27 years and in seven years time, how old will each person be?

Answer by macston(5194)

.
N=Alison's age; R=Amir's age=N-5yrs
N+R=27 years Substitute for R
N+N-5yrs=27yrs
2N=32yrs
N=16yrs Alison is currently 16 years old.
R=N-5yrs=16yrs-5yrs=11yrs Amir is currently 11 years old.
In seven years, Alison will be (16yrs+7yrs) 23 years old.
In seven years, Amir will be (11yrs+7yrs) 18 years old.